OpenStudy (anonymous):

The area of a rectangle is given by A = x2 – 2x + 30. Find the value of x if the total area of the rectangle is equal to 45 square units

OpenStudy (anonymous):

area=x2-2x+30=45 x2-2x+30-45=0 x2-2x-15=0 x2-5x+3x-15=0 x(x-5)+3(x-5)=0 (x-5)(x+3)=0 x=5 or 3